# Floor Plan-Guided Visual Navigation Incorporating Depth and Directional Cues

### v1

@@ Introduction
To address this limitation, recent work (Li et al. 2021; Ewe et al. 2024; Li et al. 2025; Setalaphruk et al. 2003; Ericson and Jensfelt 2024) has increasingly explored integrating prior floor plans with visual inputs. Floor plans constitute a valuable and widely available source of global spatial knowledge, providing high-level semantic and geometric layouts that remain invariant to dynamic changes such as human activity, temporary obstacles, or furniture rearrangement. This makes them well-suited as priors for supporting localization and navigation in unfamiliar environments. Despite their promise, existing methods still face two key challenges in fully utilizing the structural cues provided by floor plans to guide navigation effectively.
@@
First, due to the inherent modality inconsistency between RGB observations and floor plans, effectively leveraging complementary clues from these two sources to provide local obstacle avoidance and global path planning guidance remains challenging. While prior approaches bypass this challenge by establishing extra topological or semantic maps, combining them with perceptions from LiDAR (Ewe et al. 2024; Li et al. 2021) and inertial measurement units (Goswami et al. 2024), they also introduce increased system costs. FloDiff (Li et al. 2025) directly fuses latent features from floor plans and visual observations using transformer-based encoders, and performs end-to-end navigation within a diffusion policy framework. However, this straightforward fusion approach suffers from low data utilization efficiency and fails to exploit the critical geometric clues inherent in the inputs effectively.
@@
Second, prior work (Li et al. 2021, 2025) has shown that the accuracy of agent localization on the floor plan significantly affects navigation performance. Meanwhile, the spatial geometry misalignment between visual observations and floor plans makes achieving precise localization particularly challenging. Existing studies (Chen et al. 2024; Boniardi et al. 2019; Howard-Jenkins, Ruiz-Sarmiento, and Prisacariu 2021) tackle this challenge by matching geometric features from visual observations with boundary features of floor plans. However, these methods often struggle in environments with complex furniture arrangements or repetitive layouts, leading to inaccurate localization.
@@
To address the above challenges, we propose GlocDiff, a diffusion-based policy for floor plan-guided visual navigation.The core idea of GlocDiff lies in leveraging depth and directional cues from RGB observations and the floor plan, respectively, to guide action generation. As shown in Fig. 1, GlocDiff compute the shortest path from the agent’s current location to the goal on the floor plan, which—despite potential intersections with obstacles—serves as a corse yet informative directional prior. Additionally, GlocDiff leverages a pre-trained model (Ke et al. 2024) to produce depth features from RGB observations, enabling a better understanding of local scene geometry and identifying traversable regions, which enables safe navigation in complex environments. The path and depth guidance explicitly exploit geometric clues essential for next-step planning, thus bridging the modality gap inherent in the inputs and improving data efficiency.
@@
Given the reliance of navigation on high-precision localization, we further improve GlocDiff in two ways. (i) Unlike previous studies, we employ a standard visual feature-based VO algorithm (Luigi 2025) as the localization module, ensuring more stable and reliable localization throughout the navigation process. (ii) We introduce noise into the current pose and path planning results during training to enhance the model’s robustness to localization inaccuracies.
@@
Extensive experiments on the FloNa (Li et al. 2025) benchmark demonstrate that our approach achieves superior performance even with fewer training pairs, highlighting both its effectiveness and data efficiency. We also conduct comprehensive ablation studies to assess the impact of conditioning on the global shortest path and local depth features, as well as GlocDiff’s robustness to varying localization accuracy. Finally, we validate the practical applicability of GlocDiff by deploying it on an Automated Guided Vehicle (AGV) without any fine-tuning, showcasing its potential for real-world deployment in embodied systems.
@@
Mapping-based navigation, the prevailing framework, requires building an environmental map first to support subsequent planning and localization. Typically, VSLAM techniques (Mur-Artal, Montiel, and Tardós 2015; Mur-Artal and Tardós 2017) rely on visual observation inputs to construct sparse metric maps of the surrounding environment. In recent years, more researchers have integrated semantic information into map representations. Previous studies (Chaplot et al. 2020a; Ramakrishnan et al. 2022; Guo et al. 2024; Yu et al. 2024) extract semantic categories from observations, associate the categories with each point in the point cloud, and finally generate a semantic map via differentiable projection operations and a denoising network. Other approaches integrate point-level features (Huang et al. 2023), image features (Pan et al. 2025; Chen et al. 2022; Wang et al. 2023), and panoramic images (Chaplot et al. 2020b) into the built topological graph. VLFM (Yokoyama et al. 2024) iteratively constructs an occupancy map and a value map to guide the agent toward the out-of-view target object. However, the mapping-based method has to explore unseen areas, which tends to be time-consuming for navigation. Moreover, the dynamic nature of objects in indoor environments necessitates real-time updates to the constructed map, making its reuse difficult and significantly reducing efficiency.
@@
Without relying on mapping, mapless visual navigation methods directly focus on translating observations into actions. Wijmans et al. (2020) solve the point goal navigation with a near-perfect success rate under idealized settings through large-scale reinforcement learning. Learning-based visual odometry (Partsey et al. 2022; Zhao et al. 2021) is used to address realistic point goal navigation. Wenzel et al.Recent work also leverages LLMs to guide agents in navigation (Zheng et al. 2024; Chen et al. 2025; Zhou, Hong, and Wu 2024). Leveraging the diffusion policy (Chi et al. 2023), NoMaD (Sridhar et al. 2024) produces collision-free actions and flexibly handles goal-conditioned and goal-agnostic navigation via the goal masking technique. Although mapless methods eliminate the need for mapping, they often make myopic decisions due to the lack of global navigable information.
@@
Compared to the two methods mentioned above, floor plan-based navigation eliminates the need for active exploration and instead utilizes the inherent structure and spatial information to perform reliable navigation actions. However, floor plans typically only depict permanent structures such as walls and lack information about obstacles like furniture or other objects. This absence poses challenges for collision-free planning. Setalaphruk et al. (Setalaphruk et al. 2003) extract a Voronoi diagram from the sketch map and subsequently implement localization and navigation based on multiple hypotheses. Li et al. (2021) propose a cognitive navigation solution using floor plans. However, its dependency on pre-collected visual features hinders its practical applicability. To eliminate the dependency, Ewe et al. (Ewe et al. 2024) develop a graph-based navigation framework, enabling navigation without relying on precise scale information. Goswami et al. (2024) leverage stereo-inertial sensors to construct a 2D semantic point cloud, which, in conjunction with the floor plan, enables robot localization and subsequent navigation. These methods all utilize sensors other than monocular cameras. The closest related work to ours is FloNa (Li et al. 2025), which marks the first attempt to navigate unseen environments using only a monocular camera and a floor plan. FloNa utilizes a transformer-based vision encoder to extract observation context from the floor plan and visual inputs. A diffusion policy (Chi et al. 2023), conditioned on the context, is then employed to generate actions. Although FloNa demonstrates impressive performance, it fails to effectively integrate complementary geometric clues from observations and floor plans. GlocDiff tackles this issue by independently processing global and local information to derive two distinct types of guidance.
@@
Diffusion-based policies have been successfully applied in various robotic domains, including dexterous manipulation (Simeonov et al. 2023; Chi et al. 2023), locomotion (Huang et al. 2024), and path planning (Sridhar et al. 2024; Li et al. 2025). By framing robot behavior generation as a conditional denoising diffusion process, these policies not only offer high training stability but also ensure temporal consistency by generating coherent action sequences as a whole. Building upon these advances, our work focuses on utilizing diffusion policies conditioned on geometric clues derived from both visual observations and floor plans to achieve vision-based floor plan navigation.
@@
Specifically, we employ a depth encoder (Ke et al. 2024) to extract depth features from RGB images. This encoder utilizes a latent diffusion model to model the conditional distribution $P(\mathbf{d}|\mathbf{o})$, where $\mathbf{d}$ and $\mathbf{o}$ denote the low-dimensional depth feature and RGB observation, respectively. In our model, the depth encoder takes $5$ observations as input, including the current observation and the past $4$ observations, denoted as $\mathcal{O}=\{o_{t-4},o_{t-3},\ldots,o_{t}\}$. Then, the encoder outputs corresponding depth features denoted as $\mathcal{D}=\{d_{t-4},d_{t-3},\ldots,d_{t}\}$. These features are then passed through the multi-head attention module (Waswani et al. 2017), and the resulting fused features are averaged to derive the depth context. Subsequently, following (Li et al. 2025), we employ a floor plan encoder consisting of an EfficientNet-B0 (Tan and Le 2019) network and a multilayer perceptron (MLP) to encode the floor plan image, obtaining a floor plan feature. Finally, the depth context, the floor plan feature, and the agent’s pose on the floor plan from the localization module are then concatenated and fed into an MLP to produce the local depth cue, $C_{l}$.
@@
During training, we use the ground truth (GT) poses from FloNa (Li et al. 2025) as the localization results. For testing and deployment, we utilize a visual odometry or a floor plan-based visual localization module to estimate the current pose.
@@
Diffusion policy (Chi et al. 2023) leverages the Denoising Diffusion Probabilistic Models (DDPMs) (Ho, Jain, and Abbeel 2020) to formulate visuomotor robot policies. In our work, using the depth and directional cues, $C_{l}$ and $C_{g}$, the diffusion policy learns to approximate the conditional distribution $P(A_{t}|C_{l},C_{g})$ through a forward and reverse process, where $A_{t}$ is a $32\times 2$ vector representing the generated sequence of future $32$ actions at time step $t$. Each action corresponds to an approximate movement of $4$ cm, and the agent executes the first 20 steps (i.e., around $0.8$ cm) of the predicted sequence during deployment.
@@
In the reverse process, diffusion policy learns the conditional distribution $P(A_{t}^{k-1}|A_{t}^{k},C_{l},C_{g})$. Specifically, the diffusion policy employs a noise prediction network of UNet (Ronneberger, Fischer, and Brox 2015) architecture to predict the added noise, $\epsilon^{\prime}$, and utilizes the noise scheduler to carry out the denoising operation. The process can be described as:
@@
We use the data provided by FloNa (Li et al. 2025), collected from $67$ Gibson simulation environments. Each scene contains $150$, $180$, or $200$ episodes, depending on the scene size, resulting in a total of $20,214$ episodes. We sample training data from this dataset, with each sample consisting of five consecutive frames, the current and goal positions, and a floor plan image.
@@
In our implementation, the multi-head attention consists of four heads, each with four layers. We choose Square Cosine Noise Schedule (Nichol and Dhariwal 2021) to train the diffusion policy. We use the AdamW optimizer, with the learning rate controlled by PyTorch’s CosineAnnealingLR scheduler. The maximum learning rate is set to $0.0001$, and the maximum number of iterations is $20$. We set $\alpha$ to $0.1$. The training of GlocDiff is performed on $8$ NVIDIA $4090$ GPUs, with a batch size of $64$ per GPU. GlocDiff is trained for $N=20$ epochs and reaches convergence.
@@
Loc-FloDiff (F3) (Li et al. 2025) F3Loc (Chen et al. 2024) Loc-FloDiff 180180M
Loc-FloDiff (GT) (Li et al. 2025) GT Pose Loc-FloDiff 180180M
@@
We conduct our experiments in the Gibson (Xia et al. 2018) simulator. The testing episodes are sourced from FloNa (Li et al. 2025), comprising fifty unseen scenes, with each scene containing ten episodes. To ensure safe navigation, we check whether the proposed action would cause a collision before execution. If so, we intervene by instructing the agent to perform a 45° left turn and re-plan a new action sequence. The navigation is deemed successful if the final agent position falls within $\tau_{d}$ meters of the goal and the total number of interventions does not exceed $\tau_{c}$.
In our experiments, we select six baselines, with their specific components outlined in Tab. 1. Both Loc-FloDiff (F3) and GlocDiff (F3) employ F3Loc (Chen et al. 2024), an SOTA end-to-end floor plan localization model, as their localization module. Both Loc-FloDiff (GT) and GlocDiff (GT) use the ground truth (GT) pose as localization results. The Loc-A (GT) leverages ground-truth localization and directly follows the shortest path computed by the A algorithm on the floor plan. In GlocDiff (VO), a VO algorithm (Luigi 2025) is used to track the agent’s pose, serving as the localization module. Due to the lack of scale information, we manually set the scale to 0.040.04m, based on the average inter-frame displacement in our dataset.
@@ Appendix B Performance of VO and F3Loc
We additionally evaluate the accuracy of two localization modules, i.e. VO (Luigi 2025) and F3Loc (Chen et al. 2024). The input images for localization are of size $512\times 512\times 3$, with both the vertical and horizontal fields of view set to $90$ degrees. We employ the multi-frame inference mode of the pre-trained F3Loc model to estimate camera poses. For VO, we initialize the camera pose of the first frame using the ground truth. Due to the lack of an absolute scale, we manually specify the initial displacement between the first and second frames as $0.04$ m, corresponding to the average inter-frame motion observed in our dataset. We randomly sample a total of fifty episodes from ten scenes and compute the mean and median errors between the predicted and ground-truth poses at each time step. As shown in Tab. A1, VO demonstrates a higher average accuracy compared to F3Loc. Moreover, F3Loc exhibits occasional outliers, suggesting that its estimated poses abruptly jump to distant locations within the environment. Such discontinuous pose estimations are intolerable for GlocDiff. In contrast, although VO is susceptible to cumulative drift, the noise introduced remains within a range that GlocDiff can robustly tolerate, making VO a more suitable choice in practice.

### v2

@@ Introduction
To address this limitation, recent work (li2021cognitive; ewe2024spatial; li2024flona; setalaphruk2003robot; ericson2024beyond) has increasingly explored integrating prior floor plans with visual inputs. Floor plans constitute a valuable and widely available source of global spatial knowledge, providing high-level semantic and geometric layouts that remain invariant to dynamic changes such as human activity, temporary obstacles, or furniture rearrangement. This makes them well-suited as priors for supporting localization and navigation in unfamiliar environments. Despite their promise, existing methods still face two key challenges in fully utilizing the structural cues provided by floor plans to guide navigation effectively.
@@
First, due to the inherent modality inconsistency between RGB observations and floor plans, effectively leveraging complementary clues from these two sources to provide local obstacle avoidance and global path planning guidance remains challenging. While prior approaches bypass this challenge by establishing extra topological or semantic maps, combining them with perceptions from LiDAR (ewe2024spatial; li2021cognitive) and inertial measurement units (goswami2024floor), they also introduce increased system costs. FloDiff (li2024flona) directly fuses latent features from floor plans and visual observations using transformer-based encoders, and performs end-to-end navigation within a diffusion policy framework. However, this straightforward fusion approach suffers from low data utilization efficiency and fails to exploit the critical geometric clues inherent in the inputs effectively.
@@
Second, prior work (li2021cognitive; li2024flona) has shown that the accuracy of agent localization on the floor plan significantly affects navigation performance. Meanwhile, the spatial geometry misalignment between visual observations and floor plans makes achieving precise localization particularly challenging. Existing studies (chen2024f3loc; boniardi2019robot; howard2021lalaloc) tackle this challenge by matching geometric features from visual observations with boundary features of floor plans. However, these methods often struggle in environments with complex furniture arrangements or repetitive layouts, leading to inaccurate localization.
@@
To address the above challenges, we propose GlocDiff, a diffusion-based policy for floor plan-guided visual navigation.The core idea of GlocDiff lies in leveraging depth and directional cues from RGB observations and the floor plan, respectively, to guide action generation. As shown in Fig. 1, GlocDiff compute the shortest path from the agent’s current location to the goal on the floor plan, which—despite potential intersections with obstacles—serves as a corse yet informative directional prior. Additionally, GlocDiff leverages a pre-trained model (ke2024repurposing) to produce depth features from RGB observations, enabling a better understanding of local scene geometry and identifying traversable regions, which enables safe navigation in complex environments. The path and depth guidance explicitly exploit geometric clues essential for next-step planning, thus bridging the modality gap inherent in the inputs and improving data efficiency.
@@
Given the reliance of navigation on high-precision localization, we further improve GlocDiff in two ways. (i) Unlike previous studies, we employ a standard visual feature-based VO algorithm (freda2025pySLAM) as the localization module, ensuring more stable and reliable localization throughout the navigation process. (ii) We introduce noise into the current pose and path planning results during training to enhance the model’s robustness to localization inaccuracies.
@@
Extensive experiments on the FloNa (li2024flona) benchmark demonstrate that our approach achieves superior performance even with fewer training pairs, highlighting both its effectiveness and data efficiency. We also conduct comprehensive ablation studies to assess the impact of conditioning on the global shortest path and local depth features, as well as GlocDiff’s robustness to varying localization accuracy. Finally, we validate the practical applicability of GlocDiff by deploying it on an Automated Guided Vehicle (AGV) without any fine-tuning, showcasing its potential for real-world deployment in embodied systems.
@@
Mapping-based navigation, the prevailing framework, requires building an environmental map first to support subsequent planning and localization. Typically, VSLAM techniques (mur2015orb; mur2017orb) rely on visual observation inputs to construct sparse metric maps of the surrounding environment. In recent years, more researchers have integrated semantic information into map representations. Previous studies (cha2020obj; ram2022pon; guo2024object; yu2024trajectory) extract semantic categories from observations, associate the categories with each point in the point cloud, and finally generate a semantic map via differentiable projection operations and a denoising network. Other approaches integrate point-level features (huang2023visual), image features (pan2025planning; chen2022think; wang2023dreamwalker), and panoramic images (chaplot2020neural) into the built topological graph. VLFM (yokoyama2024vlfm) iteratively constructs an occupancy map and a value map to guide the agent toward the out-of-view target object. However, the mapping-based method has to explore unseen areas, which tends to be time-consuming for navigation. Moreover, the dynamic nature of objects in indoor environments necessitates real-time updates to the constructed map, making its reuse difficult and significantly reducing efficiency.
@@
Without relying on mapping, mapless visual navigation methods directly focus on translating observations into actions. wijmans2020ddppo solve the point goal navigation with a near-perfect success rate under idealized settings through large-scale reinforcement learning. Learning-based visual odometry (partsey2022ismapping; zhao2021surprising) is used to address realistic point goal navigation. Recent work also leverages LLMs to guide agents in navigation (Zheng_2024_CVPR; chen2025affordances; zhou2024navgpt). Leveraging the diffusion policy (chi2023diffusion), NoMaD (sridhar2024nomad) produces collision-free actions and flexibly handles goal-conditioned and goal-agnostic navigation via the goal masking technique. Although mapless methods eliminate the need for mapping, they often make myopic decisions due to the lack of global navigable information.
@@
Compared to the two methods mentioned above, floor plan-based navigation eliminates the need for active exploration and instead utilizes the inherent structure and spatial information to perform reliable navigation actions. However, floor plans typically only depict permanent structures such as walls and lack information about obstacles like furniture or other objects. This absence poses challenges for collision-free planning. Setalaphruk et al. (setalaphruk2003robot) extract a Voronoi diagram from the sketch map and subsequently implement localization and navigation based on multiple hypotheses. li2021cognitive propose a cognitive navigation solution using floor plans. However, its dependency on pre-collected visual features hinders its practical applicability. To eliminate the dependency, Ewe et al. (ewe2024spatial) develop a graph-based navigation framework, enabling navigation without relying on precise scale information. goswami2024floor leverage stereo-inertial sensors to construct a 2D semantic point cloud, which, in conjunction with the floor plan, enables robot localization and subsequent navigation. These methods all utilize sensors other than monocular cameras. The closest related work to ours is FloNa (li2024flona), which marks the first attempt to navigate unseen environments using only a monocular camera and a floor plan. FloNa utilizes a transformer-based vision encoder to extract observation context from the floor plan and visual inputs. A diffusion policy (chi2023diffusion), conditioned on the context, is then employed to generate actions. Although FloNa demonstrates impressive performance, it fails to effectively integrate complementary geometric clues from observations and floor plans. GlocDiff tackles this issue by independently processing global and local information to derive two distinct types of guidance.
@@
Diffusion-based policies have been successfully applied in various robotic domains, including dexterous manipulation (simeonov2023rpdiff; chi2023diffusion), locomotion (huang2024diffuseloco), and path planning (sridhar2024nomad; li2024flona). By framing robot behavior generation as a conditional denoising diffusion process, these policies not only offer high training stability but also ensure temporal consistency by generating coherent action sequences as a whole. Building upon these advances, our work focuses on utilizing diffusion policies conditioned on geometric clues derived from both visual observations and floor plans to achieve vision-based floor plan navigation.
@@
Specifically, we employ a depth encoder (ke2024repurposing) to extract depth features from RGB images. This encoder utilizes a latent diffusion model to model the conditional distribution $P(\mathbf{d}|\mathbf{o})$, where $\mathbf{d}$ and $\mathbf{o}$ denote the low-dimensional depth feature and RGB observation, respectively. In our model, the depth encoder takes $5$ observations as input, including the current observation and the past $4$ observations, denoted as $\mathcal{O}=\{o_{t-4},o_{t-3},\ldots,o_{t}\}$. Then, the encoder outputs corresponding depth features denoted as $\mathcal{D}=\{d_{t-4},d_{t-3},\ldots,d_{t}\}$. These features are then passed through the multi-head attention module (vaswani2017attention), and the resulting fused features are averaged to derive the depth context. Subsequently, following (li2024flona), we employ a floor plan encoder consisting of an EfficientNet-B0 (tan2019efficientnet) network and a multilayer perceptron (MLP) to encode the floor plan image, obtaining a floor plan feature. Finally, the depth context, the floor plan feature, and the agent’s pose on the floor plan from the localization module are then concatenated and fed into an MLP to produce the local depth cue, $C_{l}$.
@@
During training, we use the ground truth (GT) poses from FloNa (li2024flona) as the localization results. For testing and deployment, we utilize a visual odometry or a floor plan-based visual localization module to estimate the current pose.
@@
Diffusion policy (chi2023diffusion) leverages the Denoising Diffusion Probabilistic Models (DDPMs) (ho2020denoising) to formulate visuomotor robot policies. In our work, using the depth and directional cues, $C_{l}$ and $C_{g}$, the diffusion policy learns to approximate the conditional distribution $P(A_{t}|C_{l},C_{g})$ through a forward and reverse process, where $A_{t}$ is a $32\times 2$ vector representing the generated sequence of future $32$ actions at time step $t$. Each action corresponds to an approximate movement of $4$ cm, and the agent executes the first 20 steps (i.e., around $0.8$ cm) of the predicted sequence during deployment.
@@
In the reverse process, diffusion policy learns the conditional distribution $P(A_{t}^{k-1}|A_{t}^{k},C_{l},C_{g})$. Specifically, the diffusion policy employs a noise prediction network of UNet (ronneberger2015unet) architecture to predict the added noise, $\epsilon^{\prime}$, and utilizes the noise scheduler to carry out the denoising operation. The process can be described as:
@@
We use the data provided by FloNa (li2024flona), collected from $67$ Gibson simulation environments. Each scene contains $150$, $180$, or $200$ episodes, depending on the scene size, resulting in a total of $20,214$ episodes. We sample training data from this dataset, with each sample consisting of five consecutive frames, the current and goal positions, and a floor plan image.
@@
In our implementation, the multi-head attention consists of four heads, each with four layers. We choose Square Cosine Noise Schedule (nichol2021improved) to train the diffusion policy. We use the AdamW optimizer, with the learning rate controlled by PyTorch’s CosineAnnealingLR scheduler. The maximum learning rate is set to $0.0001$, and the maximum number of iterations is $20$. We set $\alpha$ to $0.1$. The training of GlocDiff is performed on $8$ NVIDIA $4090$ GPUs, with a batch size of $64$ per GPU. GlocDiff is trained for $N=20$ epochs and reaches convergence.
@@
Loc-FloDiff (F3) (li2024flona) F3Loc (chen2024f3loc) Loc-FloDiff 180180M
Loc-FloDiff (GT) (li2024flona) GT Pose Loc-FloDiff 180180M
@@
We conduct our experiments in the Gibson (xia2018gibson) simulator. The testing episodes are sourced from FloNa (li2024flona), comprising fifty unseen scenes, with each scene containing ten episodes. To ensure safe navigation, we check whether the proposed action would cause a collision before execution. If so, we intervene by instructing the agent to perform a 45° left turn and re-plan a new action sequence. The navigation is deemed successful if the final agent position falls within $\tau_{d}$ meters of the goal and the total number of interventions does not exceed $\tau_{c}$.
In our experiments, we select six baselines, with their specific components outlined in Tab. 1. Both Loc-FloDiff (F3) and GlocDiff (F3) employ F3Loc (chen2024f3loc), an SOTA end-to-end floor plan localization model, as their localization module. Both Loc-FloDiff (GT) and GlocDiff (GT) use the ground truth (GT) pose as localization results. The Loc-A (GT) leverages ground-truth localization and directly follows the shortest path computed by the A algorithm on the floor plan. In GlocDiff (VO), a VO algorithm (freda2025pySLAM) is used to track the agent’s pose, serving as the localization module. Due to the lack of scale information, we manually set the scale to 0.040.04m, based on the average inter-frame displacement in our dataset.
@@ Conclusion
Figure A1: Failure case. The left column presents the GT path and the generated path on the floor plan. The right column showcases an RGB observation along with the predicted depth and visualizes the GT path, the generated path, and the shortest path on the observation.
Table A1: Error Comparison of Localization Methods
localization method Translation (m) Orientation ()
mean median max mean median max
VO $1.53$ $1.66$ $4.28$ $13.35$ $18.72$ $22.51$
F3Loc $3.44$ $2.17$ $9.47$ $18.14$ $17.96$ $30.64$
@@ Appendix B Performance of VO and F3Loc
We additionally evaluate the accuracy of two localization modules, i.e. VO (freda2025pySLAM) and F3Loc (chen2024f3loc). The input images for localization are of size $512\times 512\times 3$, with both the vertical and horizontal fields of view set to $90$ degrees. We employ the multi-frame inference mode of the pre-trained F3Loc model to estimate camera poses. For VO, we initialize the camera pose of the first frame using the ground truth. Due to the lack of an absolute scale, we manually specify the initial displacement between the first and second frames as $0.04$ m, corresponding to the average inter-frame motion observed in our dataset. We randomly sample a total of fifty episodes from ten scenes and compute the mean and median errors between the predicted and ground-truth poses at each time step. As shown in Tab. A1, VO demonstrates a higher average accuracy compared to F3Loc. Moreover, F3Loc exhibits occasional outliers, suggesting that its estimated poses abruptly jump to distant locations within the environment. Such discontinuous pose estimations are intolerable for GlocDiff. In contrast, although VO is susceptible to cumulative drift, the noise introduced remains within a range that GlocDiff can robustly tolerate, making VO a more suitable choice in practice.